OpenStudy (anonymous):

Billy earns $25,000 the first year and gets a 5% raise each year. Write a possible explicit rule to describe its pattern. How much will he earn per year after 5 years? 10 years?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

He earns $25,000. That means $25,000 is 100% of his salary, right?

OpenStudy (anonymous):

Well that is his starting salary the first year

OpenStudy (mathstudent55):

If he gets a 5% raise, the first year of the raise he will receive 105% of his salary.

OpenStudy (anonymous):

5% of 25000 is equal to 1250

OpenStudy (anonymous):

but yes $25,000 is 100% of his starting salary the first year

OpenStudy (mathstudent55):

105% = 1.05

OpenStudy (anonymous):

1250 * 5 years

OpenStudy (anonymous):

and then add to 25000

OpenStudy (anonymous):

Y'all are confusing me

OpenStudy (anonymous):

@yahya90 I multiply 1250 by 5 then add 25000 to get my answer?

OpenStudy (mathstudent55):

Year Salary 1 25,000 2 1.05 * 25,000 3 1.05 * (1.05 * 25,000) 4 1.05 * [1.05(1.05 * 25,000)] 5 1.05 * {1.05 * [1.05 * (1.05 * 25,000)]}

OpenStudy (anonymous):

yes for 5 years >>>for ten years 10*1250 and then add to 25000

OpenStudy (mathstudent55):

Since each year he gets a 5% raise over the previous year's salary, every year you multiply the previous year's salary by 1.05.

OpenStudy (anonymous):

Great thanks you two

OpenStudy (anonymous):

Adios

OpenStudy (mathstudent55):

No, wait. @yahya90 is wrong.

OpenStudy (anonymous):

Smashing! What is the correct answer then mate?

OpenStudy (mathstudent55):

At the fifth year the pattern is: 5 1.05 * {1.05 * [1.05 * (1.05 * 25,000)]} Salary \(= 1.05^4 * 25,000\) For year n, Salary \(= 1.05^{n - 1} * 25,000\) For each year you need to figure his salary, multiply 1.05 raised to one less than the year and multiply by 25,000.

OpenStudy (anonymous):

5% is 1250 for the first year only >>>it will change in the second year >>>may be

OpenStudy (mathstudent55):

@yahya90 You are solving this as a simple interest type of problem. The way you are treating this is that he gets 5% of his original 25,000 as a raise each year. This is a compund interest type problem. The first year of his raise, he gets 5% of 25,000 = $26,250. The secnd year of the raise, he gets 5% 0f $26,250 added, not 5% of the original $25,000. etc.

OpenStudy (anonymous):

yes you are right >>>thank you

OpenStudy (mathstudent55):

wlcm